A condenser of capacity 1 μF, an inductance of 0.2 H and a resistance of 800 Ω are joined in series. Is
the circuit oscillatory? What is the natural frequency?
1
Expert's answer
2020-08-02T15:12:57-0400

As per the question,

Capacity of the condenser "c=1\mu F"

Inductance "(L)=0.2H"

Resistance "(R)=800\Omega"

For the resonating frequency,

"X_L=X_c"

"\Rightarrow \omega=\frac{1}{\sqrt{LC}}"

Now, substituting the values of L and C,

"\Rightarrow 2\pi f = \frac{1}{1\times 10^{-6}\times 0.2} sec^{-1}"


"\Rightarrow f=\frac{10^{7}}{2\times 2\pi } sec^{-1}"

"\Rightarrow f= 0.08\times 10^{7}sec^{-1}"

"\Rightarrow f=0.8\times 10^{6}sec^{-1}"